OpenStudy (anonymous):

zero exponent simplify. leave the answer in exponential notation. 1) 4^3•4^0•4^5 2) 5x^4•6x^3 3) (-3y^5)(-7y^4) 4) 32x^9y^4/24x^5y^3 5) (4x+5)^0 6) 4x^0+5 7) 8x^0+1 8) 49a^4bc^10/-7abc^8 evaluate each of the following for x=3. 9) -x^0 10) (-x)^0 11) (5x)^0 12) 5x^0 PLEASE HELP ME WITH THIS ,..

OpenStudy (anonymous):

@ganeshie8 PLEASE help me with this ..

OpenStudy (unklerhaukus):

\[x^n\cdot x^m=x^{n+m}\] \[4^3\cdot4^0\cdot4^5=4^{3+0+5}=...\]

OpenStudy (unklerhaukus):

can you do the first one?

OpenStudy (anonymous):

ok i can do .. what about the second ?

OpenStudy (anonymous):

Your constant is 5 and 6 so you multiply them add your x and then add your exponets

OpenStudy (unklerhaukus):

\[ax^n\cdot b x^m= (ab)\cdot x^{n+m}\]